# Antithetic Noise in Diffusion Models

### v1

@@ A.3.1 Monte Carlo estimator:
[…]
@@
[…]
@@
[…]
@@
Mean-squared error.
@@
A.3.2 Antithetic Monte Carlo Estimator:
@@
where we have written $\operatorname{Cov}(S_{i}^{+},S_{i}^{-})=\rho\,\sigma^{2}$ as in the statement. Set $\hat{\mu}_{N}^{\text{AMC}}=K^{-1}\sum_{i=1}^{K}\bar{S}_{i}$ and let $(\hat{\sigma}_{N}^{\text{AMC}})^{2}$ be the sample variance of $\bar{S}_{1},\dots,\bar{S}_{K}$. All three claims in Proposition 4 follow from Proposition  2.
@@ Appendix B Additional Experiments
B.1 Additional experiments on Pixel-Wise Correlations
B.1.1 DDPM
@@
Table 5 reports Pearson correlations between pixel values of image pairs produced using the same pre-trained models under PN and RR noise schemes using DDPM (default 1000 steps) across CIFAR10, CelebA-HQ, and LSUN-Church. For each dataset, we follow the same setup explained in Section 3. We calculate the pixelwise correction with 1600 antithetic noise pairs and 1600 independent noises.
With the same behavior in DDIM, in DDPM sampling, PN pairs also consistently exhibit lower correlation than RR pairs. The mean correlation for PN pairs is strongly negative in all datasets—CIFAR10 ($-0.73$), Church ($-0.45$), and in the more identity-consistent CelebA ($-0.18$). In addition, the 75th percentile of PN pairs remains below zero in all datasets, with a high of $-0.05$ (CelebA), meaning that even the upper quartile of PN correlations is still less than the average RR correlation.
Table 5: DDPM Pearson correlation coefficients for PN (pos–neg) and RR (rand–rand) pairs
@@
B.1.2 DPS
The same pattern persists in the DPS setting, too. Since DPS is guided using a reference image and there is a ground truth image available, the pixel-wise correlation is calculated using the difference between reconstructed images and the corresponding ground truth. Although the overall correlation values are shifted up—due to the deterministic conditioning and the posterior nature of sampling—PN pairs still show significantly lower correlations than RR pairs across all tasks (Inpainting, Super-resolution, Motion Deblurring).
The mean PN correlations range from $0.20$ to $0.27$, while RR correlations consistently lie above $0.50$. Importantly, the entire PN distribution remains consistently below the RR distribution: for every task, the 75th percentile of PN correlations is less than the 25th percentile of RR correlations.
@@ B.1.2 DPS
Statistic Inpainting Super-resolution Motion Deblur
PN RR PN RR PN RR
Mean 0.27 0.57 0.20 0.53 0.25 0.55
Min -0.14 0.18 -0.34 0.01 -0.32 -0.02
25th percentile 0.19 0.52 0.10 0.47 0.15 0.49
75th percentile 0.36 0.63 0.30 0.60 0.35 0.62
Max 0.66 0.82 0.64 0.80 0.69 0.81
@@
B.2 Additional experiments on the symmetry conjecture
@@
where $\bar{f}:=\int_{-1}^{1}f\mathrm{d}x/2$ is the average value of $f$ on $[-1,1]$.
The integral’s numerator is the squared average distance between the antithetic mean $0.5\,f(-x)+0.5\,f(x)$ and the overall mean $\bar{f}$, while the denominator is the full variance of $f$ over the interval.
The antisymmetry score is well-defined for every non-constant function $f$; it takes values in the range $[0,1]$ and represents the fraction of the original variance that is eliminated by antithetic averaging. The score attains $\text{AS}(f)=1$ exactly when the antithetic sum $f(x)+f(-x)$ is constant, that is, when $f$ is perfectly affine-antisymmetric. Conversely, $\text{AS}(f)=0$ if and only if $f(-x)=f(x)+c$ for some constant $c$, meaning $f$ is affine-symmetric.
[…]
@@
B.3 Additional experiments for uncertainty quantification
B.3.1 Uncertainty Quantification

### v2

@@ A.5.1 Monte Carlo estimator:
The first term converges to $\mathbb{E}[S_{1}^{2}]$ almost surely by the law of large numbers, the second term converges to $\mu^{2}$ almost surely as shown above. Therefore $(\hat{\sigma}_{N}^{\text{MC}})^{2}\xrightarrow{\text{a.s.}}\sigma^{2}$.
@@
$\sqrt{N}\,\frac{\hat{\mu}_{N}^{\text{MC}}-\mu}{\sigma}\;\xrightarrow{d}\;\mathbb{N}(0,1).$
@@
$\sqrt{N}\,\frac{\hat{\mu}_{N}^{\text{MC}}-\mu}{\hat{\sigma}_{N}^{\text{MC}}}\;\xrightarrow{d}\;\mathbb{N}(0,1).$
@@
Mean‐squared error.
@@
A.5.2 Antithetic Monte Carlo Estimator:
@@
where we have written $\operatorname{Cov}(S_{i}^{+},S_{i}^{-})=\rho\,\sigma^{2}$ as in the statement. Set $\hat{\mu}_{N}^{\text{AMC}}=K^{-1}\sum_{i=1}^{K}\bar{S}_{i}$ and let $(\hat{\sigma}_{N}^{\text{AMC}})^{2}$ be the sample variance of $\bar{S}_{1},\dots,\bar{S}_{K}$. All three claims in Proposition 6 follow from Proposition  5.
A.6 An alternative explanation via the FKG inequality
We also provide an expository discussion highlighting the FKG connection behind negative correlation. Consider the univariate case where a scalar Gaussian $z$ is fed through a sequence of one-dimensional linear maps and ReLU activations. Let $F$ be the resulting composite function. One can show that, regardless of the number of layers or the linear coefficients used, $\operatorname{Corr}\bigl(F(z),\,F(-z)\bigr)\,\leq\,0.$ The proof relies on the univariate FKG inequality (Fortuin et al., 1971), a well-known result in statistical physics and probability theory.
We generalize this result to higher dimensions via partial monotonicity, under which negative correlation still holds.
We first formally state and prove the claim for the univariate case in Section A.6.
Proposition 7 (Univariate case).
$\operatorname{Corr}\bigl(F(z),\,F(-z)\bigr)\;\leq\;0.$
Proof.
FKG inequality guarantees $\operatorname{Cov}_{Z\sim\mathbb{N}(0,1)}(f(Z),g(Z))\geq 0$ provided that $f,g$ are non-decreasing. Therefore, $\operatorname{Cov}_{Z\sim\mathbb{N}(0,1)}(F(Z),F(-Z))=-\operatorname{Cov}_{Z\sim\mathbb{N}(0,1)}(F(Z),-F(-Z))\leq 0$. Since
we have $\operatorname{Corr}\bigl(F(z),\,F(-z)\bigr)\leq 0$. ∎
To generalize the result to higher dimension, we define a function $f:\mathbb{R}^{m}\to\mathbb{R}$ to be partially monotone if for each coordinate $j$, holding all other inputs fixed, the map $t\;\mapsto\;f(x_{1},\dots,x_{j-1},t,x_{j+1},\dots,x_{m})$ is either non-decreasing or non-increasing. Mixed monotonicity is allowed. For example, f(x,y)=xyf(x,y)=x-y is non-decreasing in $x$ and non-increasing in $y$, yet still qualifies as partially monotone. We have the following result:
Proposition 8.
Now we prove the general case:
Proof of Proposition 8.
Let $G\coloneq S\circ\text{DM}$. For each coordinate $j\in[d]$ fix a sign
[…]
Similarly, define
[…]
$\operatorname{Cov}\!\bigl(\widetilde{G}(Z),-\widetilde{H}(Z)\bigr)\;\geq\;0~~\Longrightarrow~~\operatorname{Cov}\!\bigl(\widetilde{G}(Z),\widetilde{H}(Z)\bigr)\;\leq\;0~~\Leftrightarrow~~\operatorname{Cov}\!\bigl(G(Z),G(-Z)\bigr)\;\leq\;0.$
Therefore $\operatorname{Corr}\bigl(G(Z),G(-Z)\bigr)\;\leq\;0$, as claimed. ∎
Proposition 8 relies on checking partial-monotonicity. If $S$ is partially monotone (including any linear statistic), then the conditions of Proposition 8 are satisfied by, e.g., Neural Additive Models (Agarwal et al., 2021) and Deep Lattice Networks (You et al., 2017). Unfortunately, partial-monotonicity is in general hard to verify.
A.6.1 FKG inequality for DDIM
We first analyze the idealized DDIM process using the FKG inequality. A single step of the idealized DDIM is
$F_{t}(\mathbf{x})=a_{t}\mathbf{x}+c_{t}s_{t}(\mathbf{x}),$ (9)
where $s_{t}$ denotes the score of $p_{t}$, and $a_{t},c_{t}\geq 0$ are deterministic coefficients obtained by rearranging (2).
The idealized DDIM trajectory is given by the composition:
$\mathrm{DDIM}_{\mathrm{I}}=F_{1}\circ F_{2}\circ F_{3}\circ\cdots\circ F_{T}.$ (10)
We present two results that give sufficient conditions under which the local (one-step) DDIM update (9) and the global DDIM procedure (10) generate negative correlation.
We need the following definition:
Definition 1 (MTP2 with curvature bound).
Let $p:\mathbb{R}^{d}\to(0,\infty)$ be a $C^{2}$ probability density. We say that $p$ is multivariate totally positive of order 2 (MTP2) with curvature bound $\kappa\geq 0$ if the second partial derivatives of $\log p$ satisfy, for all $x\in\mathbb{R}^{d}$,
$\partial_{ij}^{2}\log p(x)\ \geq\ 0\quad\text{for all }i\neq j,\qquad\text{and}\qquad\partial_{ii}^{2}\log p(x)\ \geq\ -\kappa\quad\text{for all }i.$
MTP2 distributions are widely studied in statistics and probability Karlin and Rinott (1980). In particular, isotropic Gaussian distributions are MTP2.
Proposition 9 (One-step DDIM).
With all the notations as above, fix $t>0$, assume the distribution $p_{t}$ is MTP2 with curvature bound $\kappa_{t}$, and
$a_{t}\ \geq\ c_{t}\kappa_{t},$
then $F_{t}$ is coordinatewise nondecreasing: $\partial_{x_{j}}F_{t,i}(x)\geq 0$ for all $i,j$ and all $x$. Moreover, for any linear statistics $S$, we have $\operatorname{Corr}(S\circ F_{t}(Z)),S\circ F_{t}(-Z))\leq 0$.
Proof.
Differentiate componentwise:
$\frac{\partial F_{t,i}}{\partial x_{j}}(x)\;=\;a_{t}\,\delta_{ij}\;+\;c_{t}\,\partial_{ij}^{2}\log p_{t}(x).$
For $i\neq j$ this is $\geq 0$ by $c_{t}\geq 0$ and the mixed second derivative condition. For $i=j$,
$\frac{\partial F_{t,i}}{\partial x_{i}}(x)\ \geq\ a_{t}+c_{t}\,\big(-\kappa_{t}\big)\ \geq\ 0.$
Let $G=S\circ F_{t}$. Since $F_{t}$ is coordinately non-decreasing and $S$ is linear, we have that $G=S\circ F_{t}$ is partially monotone. Thus $\operatorname{Corr}(S\circ F_{t}(Z)),S\circ F_{t}(-Z))\leq 0$ by the multivariate FKG inequality and Proposition 8. ∎
The next result shows that the idealized DDIM trajectory continues to generate negative correlation, as long as the curvature bound holds uniformly over all steps.
Proposition 10 (Global DDIM chain).
With all the notations as above. Assume the distribution $p_{t}$ is MTP2 with curvature bound $\kappa_{t}$, and
$a_{t}\ \geq\ c_{t}\kappa_{t}$
for every $t$. Then $\mathrm{DDIM}_{\mathrm{I}}=F_{1}\circ F_{2}\circ F_{3}\circ\cdots\circ F_{T}$ is coordinatewise nondecreasing. Moreover, for any linear statistics $S$, we have $\operatorname{Corr}(S\circ\mathrm{DDIM}_{\mathrm{I}}(Z)),S\circ\mathrm{DDIM}_{\mathrm{I}}(-Z))\leq 0$.
Proof.
Since each $F_{t}$ is coordinatewise nondecreasing, their composition $F_{1}\circ\cdots\circ F_{T}$ is also coordinatewise nondecreasing. Therefore, the composition $S\circ\mathrm{DDIM}_{\mathrm{I}}$ is partially monotone, thus $\operatorname{Corr}(S\circ\mathrm{DDIM}_{\mathrm{I}}(Z)),S\circ\mathrm{DDIM}_{\mathrm{I}}(-Z))\leq 0$ as claimed. ∎
Remark 1.
All these results remain valid if we replace the exact score $\nabla\log p_{t}$ by the neural network $s_{\theta}$, up to a constant rescaling.
Remark 2.
We note that the condition $a_{t}\geq c_{t}\kappa_{t}$ is satisfied when $\kappa_{t}=O(1)$ and the probability flow ODE is discretized with sufficiently small step size. For the deterministic DDIM update under a variance-preserving schedule,
$a_{t}=\sqrt{\frac{\alpha_{t-1}}{\alpha_{t}}},\qquad c_{t}=\sqrt{1-\alpha_{t-1}}-\sqrt{\frac{\alpha_{t-1}}{\alpha_{t}}}\sqrt{1-\alpha_{t}}.$
Thus $a_{t}$ is of unit order, and $c_{t}$ captures the difference $|\alpha_{t-1}-\alpha_{t}|$. In the continuous-time limit with step size $\Delta t$ and a smooth function $\alpha(t)$, a Taylor expansion yields $a_{t}=1+O(\Delta t)$ and $c_{t}=O(\Delta t)$; thus $a_{t}\geq c_{t}\kappa_{t}$ holds when $\Delta t$ is sufficiently small and $\kappa_{t}$ stays bounded.
@@ Appendix B Additional Experiments
B.1 Correlation Experiment Setup on Other Models
B.1.1 Consistency Model
We study both unconditional and conditional consistency models using publicly available checkpoints provided in Hugging Face’s Diffusers library (von Platen et al., 2022): openai/diffusers-cd_imagenet64_l2, openai/diffusers-cd_cat256_l2, and openai/diffusers-cd_bedroom256_l2.
For unconditional models, we evaluate pre-trained models on LSUN-Cat and LSUN-Bedrooom  (Yu et al., 2015). For each dataset, 1,600 image pairs are generated under both PN and RR noise sampling with 1 DDIM steps.
For conditional models, we evaluate pre-trained models on ImageNet on 32 classes  (Deng et al., 2009). For each class, 100 PN and RR pairs are generated with 1 steps.
B.1.2 Generative Models Beyond Diffusion
A mentioned in Section 3, we evaluate correlation on a VAE and flow-based models. Unlike the diffusion models, which we use public pre-trained checkpoints, both of these models required explicit training before evaluation. Here we describe the training setup and generation procedure.
VAE:
We train the unconditional VAE on MNIST following the publicly available implementation provided in Francis (2022). The VAE consists of a simple convolutional encoder–decoder architecture with a Gaussian latent prior. After training, we generate 1,600 paired samples under PN and RR schemes, respectively.
Glow:
We train the class-conditional Glow model (Kingma and Dhariwal, 2018) on CIFAR-10 using the normflows library. The architecture follows a multiscale normalizing flow design and Glow blocks. After training, we generated 100 PN and 100 RR pairs per class.
B.2 Image Editing
We apply our antithetic initial noise strategy to the image editing algorithm FlowEdit (Kulikov et al., 2024), which edits images toward a target text prompt using pre-trained flow models.
In Algorithm 1 of FlowEdit, at each timestep, the algorithm samples $n_{\text{avg}}$ random noises $Z\sim\mathbb{N}(0,1)$ to create noisy versions of the source image, computes velocity differences between source and target conditions, and averages these directions to drive the editing process.
In the $n_{\text{avg}}=2$ setting, we replace the two independent random samples with antithetic noises: for each $Z\sim\mathbb{N}(0,I)$ we also use Z-Z and average the two velocity updates.
We compare on $76$ prompts provided in FlowEdit’s official GitHub repository. For each prompt, we generate 10 images using both PN and RR. All other parameters follow the repository defaults. We evaluate performance using CLIP(semantic text–image alignment; higher is better) and LPIPS (perceptual distance to the source; lower is better), which jointly measure text adherence and structure preservation.
As a result, PN improves the mean CLIP score, winning in 56.59% of all pairwise comparisons. It also reduces LPIPS, winning in 81.58% of all pairwise comparisons.
B.3 Implementation Detail
We use pretrained models from Hugging Face’s Diffusers library (von Platen et al., 2022): google/ddpm-church-256, google/ddpm-cifar10-32, and google/ddpm-celebahq-256 for unconditional diffusion; Stable Diffusion v1.5 for text-to-image; and the original repository from (Chung et al., 2023a) for guided generation. Experiments were run on eight NVIDIA L40 GPUs. The most intensive setup—Stable Diffusion—takes about five minutes to generate 100 images for a single prompt.
B.4 Additional experiments on pixel‐wise similarity
B.4.1 DDPM
@@
Table 5 and Table 6 report the standard Pearson correlations and centralized correlation between pixel values of image pairs produced using the same pre-trained models under PN and RR noise schemes using DDPM (default 1000 steps) across CIFAR10, CelebA-HQ, and LSUN-Church. For each dataset, we follow the same setup explained in Section 3. We calculate the pixelwise correction with 1600 antithetic noise pairs and 1600 independent noises.
Consistent with the behavior observed in DDIM, PN pairs in DDPM samples also exhibit negative correlation. Using the standard Pearson correlation, the mean correlation for PN pairs is strongly negative in all datasets—CIFAR10 (0.73-0.73), Church (0.45-0.45), and in the more identity-consistent CelebA (0.18-0.18).
The centralized correlation analysis further sharpens this contrast: mean PN correlations are substantially lower (CIFAR10 0.80-0.80, CelebA 0.67-0.67, Church 0.65-0.65). These results confirm again that PN noise pairs consistently introduce strong negative dependence, while RR pairs remain close to uncorrelated or weakly positive.
Table 5: DDPM standard Pearson correlation coefficients for PN and RR pairs
@@
Table 6: DDPM centralized Pearson correlation coefficients for PN and RR pairs
Statistic CIFAR10 CelebA Church
PN $R$ PN $R$ PN $R$
Mean -0.80 0.01 -0.67 -0.00 -0.65 -0.00
Min -0.96 -0.61 -0.89 -0.60 -0.93 -0.56
25th percentile -0.87 -0.15 -0.73 -0.15 -0.72 -0.12
75th percentile -0.75 0.16 -0.62 0.15 -0.60 0.11
Max -0.23 0.76 0.14 0.64 -0.32 0.72
@@
B.4.2 Diffusion Posterior Sampling (DPS)
The same pattern persists in the setting of using the diffusion model as a prior for posterior sampling, too, which has been utilized to solve various inverse problems, such as inpainting, super-resolution, and Gaussian deblurring.
Since there is a ground truth image available in the image restoration task, the standard pixel-wise correlation is calculated using the difference between reconstructed images and the corresponding ground truth, and the centralized correlation is calculated using the same definition described in 3. Although the overall standard correlation values are shifted up, due to the deterministic conditioning, the posterior nature of sampling—PN pairs still shows significantly lower correlations than RR pairs across all tasks.
For the standard Pearson correlation, the mean PN correlations range from $0.20$ to $0.27$, while RR correlations consistently lie above $0.50$. For the centralized correlation, PN correlations are strongly negative across all tasks (means around 0.72-0.72). In contrast, RR pairs remain centered near zero (mean correlations around 0.01-0.01 to 0.02-0.02).
@@ B.4.2 Diffusion Posterior Sampling (DPS)
Statistic Inpainting Gaussian Deblur Super-resolution
PN $R$ PN $R$ PN $R$
Mean 0.28 0.57 0.27 0.57 0.20 0.53
Min -0.14 0.05 -0.22 0.13 -0.34 0.01
25th percentile 0.19 0.52 0.17 0.51 0.10 0.47
75th percentile 0.36 0.63 0.36 0.63 0.30 0.60
Max 0.66 0.83 0.62 0.83 0.64 0.81
Table 8: DPS centralized correlation coefficients for PN and RR pairs
Statistic Inpainting Gaussian Blur Super-resolution
PN $R$ PN $R$ PN $R$
Mean -0.72 -0.02 -0.71 -0.01 -0.72 -0.01
Min -0.86 -0.43 -0.84 -0.36 -0.87 -0.35
25th percentile -0.76 -0.07 -0.75 -0.07 -0.76 -0.08
75th percentile -0.69 0.04 -0.67 0.05 -0.69 0.05
Max -0.14 0.47 0.01 0.48 -0.14 0.43
@@
B.4.3 Wasserstein distance
To complement the correlation-based analyses in the main text, we also evaluate similarity using the Wasserstein distance, a measure of distributional discrepancy. It quantifies the minimal “effort” required to transform one probability distribution into another, which means lower Wasserstein values indicate closer alignment between the two distributions, while higher values indicate larger differences.
To calculate Wasserstein distances, we treat each generated image pair as a sample from a distribution under the sampling scheme, PN or RR. As shown in Table 9, PN consistently exhibits larger Wasserstein distances than RR across nearly all models, and the differences are statistically significant at all $p<10^{-10}$. This implies that antithetic initial noises lead to more divergent distributions than random sampling and confirms our results from the correlation analysis.
Table 9: Wasserstein Distance, shown are means (SD) with corresponding t-statistics and p-values.
Model Wasserstein Distance t-stats (p)
PN $R$
LSUN-Church 0.16 (0.10) 0.12 (0.07) t=12.09,p=0t=12.09,\;p=0
CIFAR-10 0.19 (0.14) 0.15 (0.09) t=8.30,p=0t=8.30,\;p=0
CelebA-HQ 0.17 (0.11) 0.12 (0.07) t=14.50,p=0t=14.50,\;p=0
SD 1.5 0.10 (0.06) 0.09 (0.04) t=33.11,p=0t=33.11,\;p=0
DiT 0.19 (0.15) 0.14 (0.12) t=14.17,p=0t=14.17,\;p=0
VAE 0.07 (0.05) 0.05 (0.03) t=13.24,p=0t=13.24,\;p=0
Glow 0.15 (0.09) 0.12 (0.08) t=7.92,p=0t=7.92,\;p=0
B.4.4 Influence of the CFG Scale on Correlation
We extend our pixel-correlation analysis on conditional diffusion models across various Classifier-Free Guidance (CFG) scales. CFG (Ho and Salimans, 2022) is a technique that strengthens conditioning in diffusion models by interpolating between conditional and unconditional score estimates. The guidance scale controls this interpolation strength, with higher values producing samples more aligned with the conditioning signal such as prompt or class.
We conducted experiments on both Stable Diffusion 1.5 (SD1.5) and DiT using CFG scales {1, 3, 5, 7, 9}. For each setting, we generated 100 PN and RR noise pairs across 25 prompts/classes, measuring both standard and centralized correlations between the image samples.
(a) Stable Diffusion
(b) DiT
Figure 7: Average pixel correlation versus CFG scale for SD1.5 and DiT. Both standard and centralized correlations are shown. Shaded regions indicate standard deviation of mean correlation across 25 prompts/classes
As show in Figure 7, for all CFG scales, negated noise continues to produce strongly negatively correlated samples. At the same time, the standard correlation for both PN and RR grows as the CFG scale increases. This can be explained as follows: larger CFG values pull samples more strongly toward the conditioning signal (prompt/class), effectively shrinking the space of plausible outputs. As samples concentrate more tightly around the target distribution, they become more similar to one another. Thus, the correlations of both PN and RR pairs increase, while PN remains much more negative than RR.
B.4.5 Partial Negation of Noise Vectors
We conducted experiments to evaluate how localized antithetic noise influences outputs of unconditional diffusion models trained on LSUN-Church and LSUN-Bedroom. For each dataset, we generated 200 image pairs by sampling $Z_{1}\sim\mathbb{N}(0,I)$ and constructing $Z_{2}$ by negating the upper half of $Z_{1}$ while leaving the lower half unchanged.
We calculate the Pearson standard correlation and the centered correlation between corresponding halves to quantify spatial correspondence induced by the antithetic manipulation. As shown in Figure 8, the top halves exhibit sharply negative correlations, and the bottom halves are highly positively correlated and visually almost identical (Figure  27,  28). This shows that the negative correlation effect acts locally in noise space and directly affects the corresponding spatial regions in the generated images.
(a) LSUN-Church
(b) LSUN-Bedroom
Figure 8: Average pixel correlation of top-half and bottom-half across generated image pairs.
B.5 Additional experiments on the symmetry conjecture
@@
where $\bar{f}:=\int_{-1}^{1}f\mathrm{d}x/2$ is the average value of $f$ on [1,1][-1,1].
The integral’s numerator is the squared average distance between the antithetic mean 0.5f(x)+0.5f(x)0.5\,f(-x)+0.5\,f(x) and the overall mean $\bar{f}$, while the denominator is the full variance of $f$ over the interval.
The antisymmetry score is well-defined for every non-constant function $f$; it takes values in the range $[0,1]$ and represents the fraction of the original variance that is eliminated by antithetic averaging. The score attains $\text{AS}(f)=1$ exactly when the antithetic sum f(x)+f(x)f(x)+f(-x) is constant, that is, when $f$ is perfectly affine-antisymmetric. Conversely, $\text{AS}(f)=0$ if and only if f(x)=f(x)+cf(-x)=f(x)+c for some constant $c$, meaning $f$ is affine-symmetric.
For each dataset we have 100 $(5\;\text{coordinates}\times 20\;\text{time steps})$ scalar functions; the summary statistics of their AS scores are listed in Table 10. Both datasets have very high AS scores: the means exceed $0.99$ and the $10\%$ quantiles are above $0.97$, indicating that antithetic pairing eliminates nearly all variance in most cases. Even the lowest scores (about $0.77$) still remove more than three-quarters of the variance.
@@
B.6 Additional experiments for uncertainty quantification
B.6.1 Uncertainty Quantification